which of the following demostrates cummutative property?
a.2x35x5=2x5x35
b.2x35x5=2+5+35
c.2x35x5=2(30x5+5x5)
d.2x35x5=(2x30+2x5)x5
2.which of the following show distributive property?
a.2x358=(2x300)x(2x50)x(2x8)
b.2x358=(2x300)+(2x50)+(2+8)
c.2x358=(2+300)x(2+58)x(2+8)
d.2x358=(2+300)+(2+50)+2+8

Answer by KMST(5397)

which of the following demostrates commutative property?
a.2x35x5=2x5x35
b.2x35x5=2+5+35
c.2x35x5=2(30x5+5x5)
d.2x35x5=(2x30+2x5)x5
Commutation means a change or a switch, as when a sentence is switched to an easier one.
The commutative property says that changing the order of the numbers multiplied (the factors) does not change the result (the product).
a.2x35x5=2x5x35 is an example of the commutative property
b is nonsense; it is not true
c and d are examples of the distributive property.

2.which of the following show distributive property?
a.2x358=(2x300)x(2x50)x(2x8)
b.2x358=(2x300)+(2x50)+(2+8)
c.2x358=(2+300)x(2+58)x(2+8)
d.2x358=(2+300)+(2+50)+2+8
The distributive property distributes the work of multiplying a number times a sum of numbers, turning it into a few (hopefully easier multiplications) followed by adding.
a. 2x358=2x(300+50+8)=(2x300)x(2x50)x(2x8) is an example of the distributive property, and is the way we multiply.
The b, c, and d choices are not true; they are just nonsense.